# On anomalous dissipation induced by transport noise

### v2

@@ 3. Anomalous dissipation of energy – Passive scalars
This section is devoted to the proof of Theorem 1.2 which will also be a guideline for the one of Theorem 1.1. The strategy used below is outlined in Subsection 1.3.
@@ 3.1. Scaling limit at a fixed diffusivity
Proposition 3.1 (Scaling limit – passive equations).
@@
Then there exists $\delta_{0}=\delta_{0}(\gamma,\mu)>0$ for which the following assertion holds. For all $\delta\in(0,\delta_{0}]$ and sequence $(\varrho_{0,n})_{n\geq 1}\subseteq H^{\delta}$ such that $\varrho_{0,n}\to\varrho_{0}$ in $H^{\delta}$, we have
(3.1) $\lim_{n\to\infty}\mathbf{P}\Big{(}\sup_{t\in[0,1]}\|\varrho_{n}^{\gamma}(t)-\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}}\geq\varepsilon\Big{)}=0\ \text{ for all }\varepsilon>0,$
where $\varrho_{n}^{\gamma}$ and $\varrho_{\mathrm{det}}^{\gamma}$ denote the unique global solution to (4.1) with $\theta^{\gamma}=\theta^{n}$ and initial data $\varrho_{0,n}$ and the unique global solution to
(3.2) $\partial_{t}\varrho_{\mathrm{det}}^{\gamma}&=(\gamma+\mu)\Delta\varrho_{\mathrm{det}}^{\gamma}&\text{ on }&\mathbb{T}^{d},\\ \varrho_{\mathrm{det}}^{\gamma}(0,\cdot)&=\varrho_{0}&\text{ on }&\mathbb{T}^{d};\end{aligned}\right.$
@@
Solutions to (3.2) are understood as in Definition 2.1 with trivial noise. In the proof of Proposition 3.1 below, we also show that, for all $r<\delta<\delta_{0}$ and $n\geq 1$,
@@
In particular, the norm in the claim (3.1) is well-defined.
As outlined in Subsections 1.3.1 and 1.3.2, the main novelty and key point in the above result is the presence of the $L^{2}$-norm in (3.1), and the key behind such improvement is the use of the stochastic Meyers’ estimates of Section A. Actually, below we will even prove that
for all $\varepsilon>0$ and $r<\delta<\delta_{0}$, where $\delta_{0}$ is as in Proposition 3.1. However, in contrast to the 2D NSEs analysed in Section 4 below, the above will not be needed.
@@
For details, see eq. (1.9) in [61] and the comments below it. In particular, the above sequence satisfies #$\#\{k\,:\,\theta_{k}^{n}\neq 0\}<\infty$ for all $n\geq 1$.
@@
To proceed further, note that, by the Itô formula (e.g., [59, Chapter 4]) and $\nabla\cdot\sigma_{k,\alpha}=0$, the following energy equality holds for the global solution $\varrho^{\gamma}$ of (1.3):
@@
Let $\gamma\in(0,1)$, $\mu>0$ and $a\in(1,\infty)$ be fixed. Fix $\varrho_{0}\in L^{2}$, and assume that $\theta^{\gamma}\in\ell^{2}$ is normalized and radially symmetric. Let $\varrho^{\gamma}$ be the strong solution to (1.3), and set
@@
Lemma 3.3 is well-known to experts, and it is a consequence of the structure of the noise described in Subsection 2.2 and of the energy estimate (3.5), see e.g., [32, Proposition 3.6] or [1, Lemma 6.3]. For brevity, we omit the proof.
@@
The argument used here is by now well-established, the reader is referred to e.g., [37, Theorem 1.4], [32, Proposition 3.7] or [1, Theorem 6.1]. We partially repeat it to highlight the fundamental role of the estimate in Lemma 3.2 with $p>2$. Moreover, we prove the stronger result in (3.3).
@@
$C([0,1];H^{s_{0}})\cap C^{s_{0}}(0,1;H^{-r_{1}})\hookrightarrow C^{\widetilde{s}}(0,1;H^{\widetilde{r}})\hookrightarrow_{{\rm comp}}C([0,1];H^{s}),$
@@
(3.9) $\sup_{n\geq 1}\mathbf{E}\Big{[}\|\varrho^{\gamma}_{n}\|_{C([0,1];H^{r_{0}})}^{2}+\|\varrho^{\gamma}_{n}\|_{C^{r_{0}}(0,1;H^{-s_{1}})}^{2}\Big{]}<\infty.$
@@
[…]
@@
To prove (3.12), arguing as either [1, Step 2, Theorem 6.1] or [32, Proposition 3.7], it follows that
@@
In the above, we mean that the couple $(f,1)$ is a local solution to (3.2), cf., Definition 2.1 and the comments below Proposition 3.1. Now, the uniqueness of solutions to (3.2) in the class $\mathcal{X}_{N_{\gamma}}$ yields (3.12). ∎
@@
where the last equality follows from the energy balance for (3.2).
In particular, if $\varrho_{0}\neq 0$ (and therefore $\varrho_{0}\not\in{\mathbb{R}}$ as $\int_{\mathbb{T}^{d}}\varrho_{0}(x)\,\mathrm{d}x=0$), then the process $\varrho_{n}^{\gamma}$ does not convergence in probability in $L^{2}(0,1;H^{1}(\mathbb{T}^{d}))$ to $\varrho_{\mathrm{det}}^{\gamma}$.
@@
Fix $\varepsilon>0$. Let $(\theta_{n})_{n\geq 1}$ be as in (3.4), and due to the comments below it, we have $\lim_{n\to\infty}\|\theta^{n}\|_{\ell^{\infty}}=0$. To conclude, it is enough to show that
(3.14) $\lim_{n\to\infty}\sup_{\varrho_{0}\in\mathcal{B}_{N,\delta}}\mathbf{P}\Big{(}\sup_{t\in[0,1]}\|\varrho^{n}(\varrho_{0})-\varrho_{\mathrm{det}}(\varrho_{0})\|_{L^{2}}\geq\varepsilon\Big{)}=0.$
where $\varrho^{n}(\varrho_{0})$ and $\varrho_{\mathrm{det}}(\varrho_{0})$ denote the solution to (1.3) with $\theta^{\gamma}=\theta^{n}$ and (3.2) both with initial data $\varrho_{0}$, respectively. We prove (3.14) by contradiction. Indeed, assume that there exists a sub-sequence $(\varrho_{0,n_{j}})_{j\geq 1}\subseteq\mathcal{B}_{N,\delta}$ such that
(3.15) $\lim_{j\to\infty}\mathbf{P}\Big{(}\sup_{t\in[0,1]}\|\varrho_{j}-\varrho_{\mathrm{det},j}\|_{L^{2}}\geq\varepsilon\Big{)}>0.$
where $\varrho_{j}\stackrel{{\scriptstyle{\rm}}}{{=}}\varrho(\varrho_{0,n_{j}})$ and $\varrho_{\mathrm{det},j}\stackrel{{\scriptstyle{\rm}}}{{=}}\varrho(\varrho_{0,n_{j}})$. Note that, for all $\widetilde{\delta}\in(0,\delta)$, there exists a (not-relabeled) sub-sequence $(\varrho_{0,n_{j}})$ such that $\varrho_{0,n_{j}}\to\varrho_{0}$ in $H^{\widetilde{\delta}}$. Now the contradiction with (3.15) follows by applying Proposition 3.1 with $\delta=\widetilde{\delta}$. Thus, (3.14) is proved. ∎
@@
We split the proof into two cases.
(1): We begin by collecting some useful facts. Let $(\gamma,N,\delta)$ be as in the statement of Theorem 1.1. In particular,
(3.16) $N^{-1}\leq\|\varrho_{0}\|_{L^{2}}\quad\text{ and }\quad\|\varrho_{0}\|_{H^{\delta}}\leq N.$
Moreover, by the mean-zero assumption on $\varrho_{0}$ and the Poincaré inequality, it follows that the unique global solution $\varrho_{\mathrm{det}}^{\gamma}$ of (3.2) satisfies
(3.17) $\|\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}}^{2}\leq e^{-\mu t}\|\varrho_{0}\|_{L^{2}}^{2}\ \text{ for all }t>0.$
Applying Corollary 3.5 with $(\mu,N,\delta)$ as above and $\gamma>0$ fixed, we obtain a family of $(\theta^{\gamma})_{\gamma\in(0,1)}\in\ell^{2}$ for which the unique global solution $\varrho^{\gamma}$ to (1.3) satisfies
(3.18) $\mathbf{P}\Big{(}\sup_{t\in[0,1]}\|\varrho^{\gamma}-\varrho_{\mathrm{det}}^{\gamma}\|_{L^{2}}\leq\frac{\varepsilon}{2N}\Big{)}>1-\varepsilon,$
where $\varrho_{\mathrm{det}}^{\gamma}$ is the unique global solution of (3.2) and $\varepsilon$ is chosen so that
(3.19) $0<\varepsilon<\Big{(}\frac{1-e^{-\mu}}{2N}\Big{)}\wedge\frac{1}{2}.$
Now we turn to the proof of (1). Let $\tau$ be the stopping time defined as
(3.20) $\tau\stackrel{{\scriptstyle{\rm def}}}{{=}}\inf\big{\{}t\in[0,1]\,:\,\|\varrho^{\gamma}(t)-\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}}\geq\tfrac{\varepsilon}{2N}\big{\}}\ \ \text{ where }\ \ \inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}1.$
By (3.18), it follows that
@@
Note that, by (3.5), (3.16) and (3.17) show that $\|\varrho^{\gamma}(t)\|_{L^{2}}\vee\|\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}}\leq N$ for all $t\in[0,1]$. In particular, by (3.20),
$\mathbf{E}\big{[}\mathbf{1}_{\{\tau\geq 1\}}\big{|}\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}-\|\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}}^{2}\big{|}\big{]}\leq 2N\,\mathbf{E}\big{[}\mathbf{1}_{\{\tau\geq 1\}}\|\varrho^{\gamma}(1)-\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}}\big{]}\leq\varepsilon.$
Combining the above with (3.21), we obtain
@@
Hence, for all $\gamma\in(0,1)$,
(3.23) $2\,\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{2}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t$ =(3.5)${{=}}\|\varrho_{0}\|_{L^{2}}^{2}-\mathbf{E}\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}$
$\,\ \stackrel{{\scriptstyle(i)}}{{\geq}}(1-\varepsilon)\|\varrho_{0}\|_{L^{2}}^{2}-\mathbf{E}[\mathbf{1}_{\{\tau\geq 1\}}\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}]$
(3.22)${{\geq}}\|\varrho_{0}\|_{L^{2}}^{2}(1-\varepsilon)(1-e^{-\mu})-\varepsilon.$
where in $(i)$ we used (3.21) and $\|\varrho^{\gamma}(1)\|_{L^{2}}\leq\|\varrho_{0}\|_{L^{2}}$ a.s. due to (3.5). Now, the claim of Theorem 1.1(1) follows from (3.16), (3.19) and the arbitrariness of $\gamma\in(0,1)$.
(2): The proof is analogous to the one of (1), provided one chooses the parameter $\varepsilon$ also depending on the rate of dissipation $\eta$. Firstly, since $\mu>-\ln(1-\eta)$, it follows that $\eta_{1}\stackrel{{\scriptstyle{\rm def}}}{{=}}1-e^{-\mu}>\eta$. Secondly, let $\eta_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}\frac{\eta+\eta_{1}}{2}\in(\eta,\eta_{1})$ and fix
$0<\varepsilon<\Big{(}1-\frac{\eta_{0}}{\eta_{1}}\Big{)}\wedge\Big{(}\frac{\eta_{0}-\eta}{N}\Big{)}.$
Then, it is clear that $\|\varrho_{0}\|_{L^{2}}^{2}(1-\varepsilon)(1-e^{-\mu})-\varepsilon\geq\eta\|\varrho_{0}\|_{L^{2}}^{2}$ as $\|\varrho_{0}\|_{L^{2}}\geq N^{-1}$, and therefore the assertion (2) of Theorem 1.2 follows from the lower bound (3.23). ∎

### v3

@@ 3. Anomalous dissipation of energy – Passive scalars
This section is devoted to the proof of Theorem 1.2, and will also serve as a guideline for that of Theorem 1.1. The strategy used is outlined in Subsection 1.3.
@@ 3.1. Scaling limit at a fixed diffusivity
Proposition 3.1 (Scaling limit – Passive scalars).
@@
Let $\delta>0$ and $\varrho_{0}\in H^{\delta}(\mathbb{T}^{d})$. Then, for all sequences $(\varrho_{0,n})_{n\geq 1}\subseteq H^{\delta}(\mathbb{T}^{d})$ such that $\varrho_{0,n}\to\varrho_{0}$ in $H^{\delta}(\mathbb{T}^{d})$, we have
(3.1) $\lim_{n\to\infty}\mathbf{P}\Big(\sup_{t\in[0,1]}\|\varrho_{n}^{\gamma}(t)-\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}(\mathbb{T}^{d})}\geq\varepsilon\Big)=0\ \text{ for all }\varepsilon>0,$
where $\varrho_{n}^{\gamma}$ and $\varrho_{\mathrm{det}}^{\gamma}$ denotes the unique global solution to (1.3) with $\theta^{\gamma}=\theta^{n}$ and initial data $\varrho_{0,n}$ and the unique global solution to
(3.2) $\left\{\begin{aligned} \partial_{t}\varrho_{\mathrm{det}}^{\gamma}&=(\gamma+\mu)\Delta\varrho_{\mathrm{det}}^{\gamma}&\text{ on }&\mathbb{T}^{d},\\ \varrho_{\mathrm{det}}^{\gamma}(0,\cdot)&=\varrho_{0}&\text{ on }&\mathbb{T}^{d};\end{aligned}\right$.
@@
Solutions to (3.2) are understood as in Definition 2.1 with trivial noise.
As outlined in Subsections 1.3.1 and 1.3.2, the main novelty and key point in the above result is the presence of the $L^{2}$-norm in (3.1), and the key behind such improvement is the use of the stochastic Meyers’ estimates of Section A. In the proof of Proposition 3.1 below, we actually prove a stronger version of Proposition 3.1. More precisely, we show the existence of $\delta_{0}=\delta_{0}(\mu,\gamma)>0$ such that, for all $\delta\in(0,\delta_{0}]$ and sequences $(\varrho_{0,n})_{n\geq 1}\subseteq H^{\delta}$ such that $\varrho_{0,n}\to\varrho_{0}$ in $H^{\delta}$, it holds that
for all $\varepsilon>0$ and $r<\delta<\delta_{0}$. We emphasize that the norm appearing in (3.1) is well-defined, since–as shown in the proof of Proposition 3.1 below–for all $r<\delta\leq\delta_{0}$ and $n\geq 1$,
@@
In contrast to the 2D NSEs analysed in Section 4 below, the stronger convergence in (3.3) will not be needed.
@@
For details, see eq. (1.9) in [67] and the comments below it. In particular, the above sequence satisfies $\{k\,:\,\theta_{k}^{n}\neq 0\}<\infty$ for all $n\geq 1$.
@@
To proceed further, note that, by the Itô formula (e.g., [65, Chapter 4]) and $\nabla\cdot\sigma_{k,\alpha}=0$, the following energy equality holds for the global solution $\varrho^{\gamma}$ of (1.3):
@@
Let $\gamma\in(0,1)$, $\mu>0$ and $a\in(1,\infty)$ be fixed. Fix $\varrho_{0}\in L^{2}$, and assume that $\theta^{\gamma}\in\ell^{2}$ is normalized and radially symmetric. Let $\varrho^{\gamma}$ be the unique global solution to (1.3), and set
@@
Lemma 3.3 is well-known to experts, and it is a consequence of the structure of the noise described in Subsection 2.2 and of the energy estimate (3.5), see e.g., [36, Proposition 3.6] or [1, Lemma 6.3]. For brevity, we omit the proof.
@@
The argument used here is by now well-established; the reader is referred to e.g., [42, Theorem 1.4], [36, Proposition 3.7] or [1, Theorem 6.1]. We partially repeat it to highlight the fundamental role of the estimate in Lemma 3.2 with $p>2$. Moreover, we prove the stronger result in (3.3).
@@
$C([0,1];H^{r_{0}})\cap C^{s_{0}}(0,1;H^{-r_{1}})\hookrightarrow C^{\widetilde{s}}(0,1;H^{\widetilde{r}})\hookrightarrow_{{\rm comp}}C([0,1];H^{r}),$
@@
(3.9) $\sup_{n\geq 1}\mathbf{E}\Big[\|\varrho^{\gamma}_{n}\|_{C([0,1];H^{r_{0}})}^{2}+\|\varrho^{\gamma}_{n}\|_{C^{s_{0}}(0,1;H^{-r_{1}})}^{2}\Big]<\infty.$
@@
(3.10) $\sup_{t\in[0,1]}\|\varrho_{n}^{\gamma}(t)\|_{L^{2}}^{2}+\int_{0}^{1}\int_{\mathbb{T}^{d}}|\nabla\varrho_{n}^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t\leq N_{\gamma}\ \text{ a.s., }$
@@
To prove (3.12), arguing as either [1, Step 2, Theorem 6.1] or [36, Proposition 3.7], it follows that
@@
In the above, we mean that the couple $(f,1)$ is a local solution to (3.2), cf. Definition 2.1 and the comments below Proposition 3.1. Now, the uniqueness and maximality of solutions to (3.2) in the class $\mathcal{X}_{N_{\gamma}}$ yields (3.12). ∎
@@
where the last equality follows from the energy balance for (3.2). In the above, we also used that $\|\varrho^{\gamma}(t)\|_{L^{2}(\mathbb{T}^{d})},\|\varrho_{\mathrm{det}}(t)\|_{L^{2}(\mathbb{T}^{d})}\leq\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}$ for all $\gamma,t>0$ a.s.
In particular, if $\varrho_{0}\neq 0$ (and therefore $\varrho_{0}\not\in{\mathbb{R}}$ as $\int_{\mathbb{T}^{d}}\varrho_{0}(x)\,\mathrm{d}x=0$), then the process $\varrho_{n}^{\gamma}$ does not convergence to $\varrho_{\mathrm{det}}^{\gamma}$ in probability in $L^{2}(0,1;H^{1}(\mathbb{T}^{d}))$.
@@
Fix $\varepsilon>0$. Let $(\theta_{n})_{n\geq 1}$ be as in (3.4), and due to the comments below it, we have $\lim_{n\to\infty}\|\theta^{n}\|_{\ell^{\infty}}=0$. To conclude, it is enough to show that, for all $\varepsilon>0$,
(3.14) $\lim_{n\to\infty}\sup_{\varrho_{0}\in\mathcal{B}_{N,\delta}}\mathbf{P}\Big(\sup_{t\in[0,1]}\|\varrho^{\gamma}_{n}(\varrho_{0})-\varrho_{\mathrm{det}}^{\gamma}(\varrho_{0})\|_{L^{2}}\geq\varepsilon\Big)=0.$
where $\varrho^{\gamma}_{n}(\varrho_{0})$ and $\varrho_{\mathrm{det}}^{\gamma}(\varrho_{0})$ denote the solution to (1.3) with $\theta^{\gamma}=\theta^{n}$ and (3.2) both with initial data $\varrho_{0}$, respectively. We prove (3.14) by contradiction. Indeed, if (3.14) is not true, then that there exists $\varepsilon_{0}>0$ and a (not-relabeled) sub-sequence $(\varrho_{0,n})_{n\geq 1}\subseteq\mathcal{B}_{N,\delta}$ such that
(3.15) $\lim_{n\to\infty}\mathbf{P}\Big(\sup_{t\in[0,1]}\|\varrho^{\gamma}_{n}(t)-\varrho_{\mathrm{det},n}^{\gamma}(t)\|_{L^{2}}\geq\varepsilon_{0}\Big)>0,$
where $\varrho^{\gamma}_{n}\stackrel{{\scriptstyle{\rm def}}}{{=}}\varrho^{\gamma}_{n}(\varrho_{0,n})$ and $\varrho_{\mathrm{det},n}^{\gamma}\stackrel{{\scriptstyle{\rm def}}}{{=}}\varrho_{\mathrm{det}}^{\gamma}(\varrho_{0,n})$. In the remaining part of this proof, we show that (3.15) leads to a contradiction with Proposition 3.1. To this end, note that, for all $\delta_{0}\in(0,\delta)$, there exists a (again, not-relabeled) sub-sequence $(\varrho_{0,n})$ such that $\varrho_{0,n}\to\varrho_{0}$ in $H^{\delta_{0}}$. From the latter, it readily follows that
(3.16) $\varrho_{\mathrm{det},n}^{\gamma}\to\varrho_{\mathrm{det}}^{\gamma}\text{ in }C([0,1];L^{2}).$
Recall from Proposition 3.1 that
(3.17) $\lim_{n\to\infty}\mathbf{P}\Big(\sup_{t\in[0,1]}\|\varrho^{\gamma}_{n}(t)-\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}}\geq\frac{\varepsilon_{0}}{2}\Big)=0.$
Hence, from the triangle inequality, (3.16) and (3.17) contradict (3.15). Thus, (3.14) is proved, and this concludes the proof. ∎
@@
We begin by collecting some useful facts. Let $N$ and $\delta$ be as in the statement of Theorem 1.1. In particular,
(3.18) $\|\varrho_{0}\|_{H^{\delta}}\leq N.$
Moreover, by the mean-zero assumption on $\varrho_{0}$ and the Poincaré inequality $\|\varrho\|_{L^{2}}\leq\frac{1}{2\pi}\|\nabla\varrho\|_{L^{2}}$, it follows that the unique global solution $\varrho_{\mathrm{det}}^{\gamma}$ of (3.2) satisfies
(3.19) $\|\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}}^{2}\leq e^{-(\mu t)/(4\pi^{2})}\|\varrho_{0}\|_{L^{2}}^{2}\ \text{ for all }t>0.$
We now split the proof into two steps.
Step 1: For all $\mu>0$ and $\gamma,\varepsilon\in(0,1)$, there exists $\theta_{\varepsilon}^{\gamma}\in\mathcal{S}_{\ell^{2}}^{0}$ such that the unique global solution to (1.3) (see Definition 2.1) satisfies
(3.20) $2\,\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq(1-\varepsilon-e^{-\mu/(4\pi^{2})})\|\varrho_{0}\|_{L^{2}}^{2}-2\varepsilon\|\varrho_{0}\|_{L^{2}}.$
Recall that $\mathcal{S}_{\ell^{2}}^{0}$ is defined in (1.4). From Corollary 3.5 applied with $\mu,N,\delta$ and $\gamma$ as above, there exists $\theta_{\varepsilon}^{\gamma}\in\mathcal{S}_{\ell^{2}}^{0}$ (depending also on the reamining parameters $\mu,N$ and $\delta$, which are not displayed as they are considered fixed) for which the unique global solution $\varrho^{\gamma}$ to (1.3) with $\theta^{\gamma}=\theta^{\gamma}_{\varepsilon}$ satisfies
where $\varrho_{\mathrm{det}}^{\gamma}$ is the unique global solution of (3.2), with $\mu>0$ as above.
Let $\tau$ be the stopping time defined as
(3.22) $\tau\stackrel{{\scriptstyle{\rm def}}}{{=}}\inf\big\{t\in[0,1]\,:\,\|\varrho^{\gamma}(t)-\varrho_{\mathrm{det}}^{\gamma}(t)\|_{L^{2}}\geq\varepsilon\big\}\ \ \text{ where }\ \ \inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}1.$
It follows from (3.21) that
@@
Applying the Itô’s formula to compute $\|\varrho(1)\|_{L^{2}}^{2}$ and using $\nabla\cdot\sigma_{k,\alpha}=0$, we have
(3.24) $2\int_{0}^{t}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t=\|\varrho_{0}\|_{L^{2}}^{2}-\|\varrho(t)\|_{L^{2}}^{2}\ \ \text{ a.s.\ for all }t>0.$
In particular,
(3.25) $2\,\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t$ $=\|\varrho_{0}\|_{L^{2}}^{2}-\mathbf{E}\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}$
$=\|\varrho_{0}\|_{L^{2}}^{2}-\mathbf{E}[\mathbf{1}_{\{\tau<1\}}\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}]-\mathbf{E}[\mathbf{1}_{\{\tau\geq 1\}}\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}].$
Next, we estimate each term separately. Firstly, as $\|\varrho^{\gamma}(1)\|_{L^{2}}\leq\|\varrho_{0}\|_{L^{2}}$ a.s., due to the energy balance (3.24), it holds that
@@
Secondly, by the definition of $\tau$, and once again invoking the energy inequalities for $\varrho^{\gamma}$ and $\varrho_{\mathrm{det}}^{\gamma}$, we have, a.s. on $\{\tau\geq 1\}$,
$\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}$ $\leq\|\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}}^{2}+\big|\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}-\|\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}}^{2}\big|$
$\leq\|\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}}^{2}+\big(\|\varrho^{\gamma}(1)\|_{L^{2}}+\|\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}}\big)\|\varrho^{\gamma}(1)-\varrho_{\mathrm{det}}^{\gamma}(1)\|_{L^{2}}$
$\leq e^{-\mu/(4\pi^{2})}\|\varrho_{0}\|_{L^{2}}^{2}+2\varepsilon\|\varrho_{0}\|_{L^{2}}.$
Therefore,
(3.27) $\mathbf{E}[\mathbf{1}_{\{\tau\geq 1\}}\|\varrho^{\gamma}(1)\|_{L^{2}}^{2}\big]\leq e^{-\mu/(4\pi^{2})}\|\varrho_{0}\|_{L^{2}}^{2}+2\varepsilon\|\varrho_{0}\|_{L^{2}}.$
Hence, (3.20) follows by putting together the estimates (3.25), (3.26) and (3.27).
Step 2: Conclusion. Let $(\varepsilon_{\gamma})_{\gamma\in(0,1)}$ be any family of positive numbers in $(0,1)$ satisfying $\lim_{\gamma\downarrow 0}\varepsilon_{\gamma}=0$. Let
(3.28) $\theta^{\gamma}\stackrel{{\scriptstyle{\rm def}}}{{=}}\theta^{\gamma}_{\varepsilon^{\gamma}}\in\mathcal{S}_{\ell^{2}}^{0}$
where $\theta^{\gamma}_{\varepsilon}$ are as in Step 1. From the inequality (3.20) and $\|\varrho_{0}\|_{L^{2}}\leq N$ by (3.18), we have
$2\liminf_{\gamma\downarrow 0}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t$ $\geq\liminf_{\gamma\downarrow 0}\big[(1-\varepsilon_{\gamma}-e^{-\mu/(4\pi^{2})})\|\varrho_{0}\|_{L^{2}}^{2}-2N\varepsilon_{\gamma}\big]$
$=(1-e^{-\mu/(4\pi^{2})})\|\varrho_{0}\|_{L^{2}}^{2},$
where $\varrho^{\gamma}$ is the unique global solution to (1.3) with $\theta^{\gamma}$ as in (3.28). ∎
As announced in Remark 1.3, the argument used to prove Theorem 1.2 yields the following slightly different version of the latter.
Proposition 3.6 (Anomalous dissipation of energy by transport noise – Passive scalars II).
Let $d\in{\mathbb{N}}_{\geq 1}$, $N\geq 1$ and $\delta>0$. For all $\mu>0$ and $\eta\in(0,1-e^{-\mu/(4\pi^{2})})$, there exists a family $(\theta^{\gamma})_{\gamma\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\varrho_{0}\in H^{\delta}(\mathbb{T}^{d})$ satisfying $N^{-1}\leq\|\varrho_{0}\|_{L^{2}}$ and $\|\varrho_{0}\|_{H^{\delta}}\leq N$, it holds that
$\inf_{\gamma\in(0,1)}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{\eta}{2}\,\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2},$
where $\varrho^{\gamma}$ is the unique global smooth solution to (1.3).
The proof requires only a minor modification of that of Theorem 1.2.
Proof.
From (3.20) in the above proof of Theorem 1.2, for each $\gamma,\varepsilon\in(0,1)$, we can find $\theta_{\varepsilon}^{\gamma}\in\mathcal{S}_{\ell^{2}}^{0}$ such that
$2\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t$ $\geq(1-\varepsilon-e^{-\mu/(4\pi^{2})})\|\varrho_{0}\|_{L^{2}}^{2}-2\varepsilon\|\varrho_{0}\|_{L^{2}}$
$\stackrel{{\scriptstyle(i)}}{{\geq}}\big[(1-\varepsilon-e^{-\mu/(4\pi^{2})})-2N\varepsilon\big]\,\|\varrho_{0}\|_{L^{2}}^{2},$
where in $(i)$ we used that $\|\varrho_{0}\|_{L^{2}}\geq N^{-1}$ by assumption. Since $\eta<1-e^{-\mu/(4\pi^{2})}$, the claimed estimate follows by choosing $\varepsilon_{\eta}$ depending only on $N$ and $\mu$ such that $\big[(1-\varepsilon_{\eta}-e^{-\mu/(4\pi^{2})})-2N\varepsilon_{\eta}\big]\geq\eta$, with a corresponding choice of $\theta^{\gamma}=\theta^{\gamma}_{\varepsilon_{\eta}}$. ∎